# BRACTIVE: A Brain Activation Approach to Human Visual Brain Learning

### v2

@@ 3.5 Subject of Interest Retrieval
As outlined in the previous section, it is infeasible to directly form $\textbf{P}_{s_{k}}$ and $\textbf{F}_{s_{k}}$. The Remark 1 and Remark 2 imply that if we can find the coefficient $\alpha_{ik}$, we can measure them directly. Fortunately, according to the Remark 3, $\mathbf{P}_{s_{k}}$, $\mathbf{F}_{s_{k}}$ and $\mathbf{T}_{s_{k}}$ are equivalent, but in the form of different modalities. Therefore, we measure the coefficient $\alpha_{ik}$ by measuring the similarity of $\mathbf{T}_{s_{k}}$ and each token. Motivated by this idea, we introduce the SOIR module to construct the features of interested subjects in visual and fMRI modalities. We detail how this module operates for visual $\textbf{P}_{s_{k}}$ and note that the approach is similarly applicable for $\textbf{F}_{s_{k}}$.
Given a sequence of visual tokens $\mathbf{P}_{v}=\left[\mathbf{P}_{0},\dots,\mathbf{P}_{i},\dots,\mathbf{P}_{N_{p}-1}\right]$. Firstly, we measure coefficient $\alpha_{ik}=\texttt{sim}(\mathbf{P}_{i},\textbf{T}_{s_{k}})$. After that, we use the softmax function over these coefficients to normalize the feature aggregation. Consequently, we form $\textbf{P}_{s_{k}}$ as the weighted sum of patches’ features as mentioned in the Section 3.2.
@@ 3.6 Objective Loss Functions
Global Contextual Loss. This loss function keeps the global context sharing between visual, text, and fMRI similar. In particular, let Contras be a contrastive loss function. The global context loss $\mathcal{L}_{g}$ is defined as in Eq. (7).
@@ 4 Brain Region of Interest Localization
As noted in Remark 2, the sub-signals or voxels in fMRI corresponding to a subject can be identified by evaluating the coefficient $\beta_{ik}$. This coefficient is defined as $\beta_{ik}=\texttt{sim}(\mathbf{F}_{s_{k}},\mathbf{F}i)$, representing the cosine similarity between $\mathbf{F}_{s_{k}}$ and $\mathbf{F}_{i}$. Additionally, a threshold $\gamma$ is predefined, designating that a sub-signal $f_{i}\in f$ is associated with the subject $s_{k}$ if $\beta_{ik}>\gamma$. The procedure for detecting the subject-specific fMRI signal is detailed in Algorithm 2. This algorithm highlights two principal observations. (1) Thanks to the Weighted SOI Loss, we can employ both kinds of features from the interested subjects, such as $\mathbf{T}_{s_{k}},\mathbf{P}_{s_{k}}$, as inputs of the algorithm. This flexibility allows for multiple approaches to be used in studying brain activities. Either text or vision modalities can be used to detect the ROIs of the subjects. (2) The Algorithm 2 can be adapted for visual analysis by substituting $\textbf{F}_{f}$ with $\mathbf{P}_{v}$. This modification allows us to identify which parts of an image strongly correlate with predetermined regions of interest (ROIs) in the brain.
@@ 5 Implementation Details
Visual Encoder and fMRI Encoder. We employ the ViT-B-16 for both visual and fMRI encoder. The input image size is $224\times 224$. The feature dimension of these encoders is set to 1024.
Text Encoder. We take the pretrained transformer architecture from [4] to extract features of the text. The maximum length of the context is set to 77 [4] and the dimension of text features is also set to 1024. The text encoder is frozen while training. It will help to align visual and fMRI features to the text features easily.
@@
Datasets. To train BRACTIVE, we use the Natural Scenes Dataset (NSD) [77], a comprehensive compilation of responses from eight participants obtained through high-quality 7T fMRI scans. Each subject was exposed to approximately 73,000 natural scenes, forming the basis for constructing visual brain encoding models. Since the visual stimulus in this database is a subset of COCO [78], for each sample, there exists captions (text modality), visual stimulus, and corresponding fMRI response, respectively. For each subject, we split the data into 5 folds for training and validation.
Training Strategy. BRACTIVE is implemented easily in Pytorch framework and trained by 16 $\times$ A100 GPUs (40G each). The learning rate is set to $2.5e^{-5}$ initially and then reduced to zero gradually under ConsineLinear [79] policy. The batch size is set to $4$/GPU. The model is optimized by AdamW [80] for $30$ epochs. The training is completed within two hours.
@@ 6 Experimental Results
Table I: The Dice Score of ROIs localization w.r.t human subjects. The AvgAtt is the average of the attention maps of human subjects across participants.
@@
Table II: Performance of the brain response prediction. IMN1K refers to pretrained weights from Imagnet1K [82]. CLS and SOI refer to the CLS token and Subject of Interest features, respectively.
@@
Since the ROIs for specific subjects are limited, in the scope of this paper, we evaluate the accuracy of BRACTIVE with respect to human subjects only. In particular, we get the ROIs of floc-faces and floc-bodies areas inside the brain as the ground truth and evaluate how well our algorithm 2 could localize the ROIs of the human subjects. We utilize dice scores as the evaluation metric. The results are demonstrated in the Table I. The details and visualization of the prediction and ground truth can be found in the supplementary section.
Figure 4: Region Of Interest of the human subject perceived by participant P1. The last column includes the mean of attention maps across all samples of all participants in the NSD. In the attention map, the brighter the colors, the higher correlation to the subject.
@@
In the second experiment, we use steps outlined in Algorithm 2 to localize the ROIs. Overall, we observe a notable enhancement ranging from 68% to 69% consistently across eight participants in comparison to the GradCam [81]. In the third experiment, we take the average of the attention maps across all participants. Surprisingly, the performance drops to 54.84 % for the GradCam [81] baseline while we observe a notable improvement of 72.57% from our method.
6.2 Brain Response Prediction
This downstream task aims to predict the neural responses of the human brain to natural scenes, as observed during participant viewing sessions [83, 66]. We employ the pretrained visual encoder $\mathcal{E}_{v}$ to fine-tune this task. Our evaluation follows the protocol described in [83], using the Pearson Correlation Coefficient (PCC) as the metric. Results, as presented in Table II, compare different architectures and training strategies. The performance for these methods on eight participants varies between 41.02 and 41.95. In a second experiment, we start with the pretrained visual encoder from BRACTIVE as initial weights and use the features from the CLS token in the final linear layer, enhancing performance by approximately 5% compared to the initial scenario. In a final experiment, settings remain similar to the second, but instead of using CLS token features, we use a weighted sum of SOI features. This adjustment leads to substantial performance improvements of 10% and 3% over the first and second experiments, respectively. The SOI features provide a richer context within the visual stimuli than the CLS features, explaining the observed enhancements in model performance.
6.3 Other Downstream Tasks: Detection and Segmentation
This section investigates how the pretrained vision encoder $\mathcal{E}_{v}$ could help on other downstream tasks. Especially, we select the most common vision tasks: object detection, instance, and semantic segmentation. For object detection and instance segmentation, we use Mask-RCNN [84] architecture, COCO [78] database for training and evaluation. For semantic segmentation, we use UpperNet [85] architecture and ADE20K [86] database. We conveniently utilize the base code from MMDetection and MMSegmentation [87].
Table III: Results of the downstream tasks: object detection, instance segmentation on COCO dataset, and semantic segmentation on ADE20K dataset
[…]
@@ 6.3 Other Downstream Tasks: Detection and Segmentation
7 Ablation Study with Qualitative Results
@@
In this section, we demonstrate the detected ROIs of human subjects as well as non-human subjects and then analyze the meaning behind them.
7.1 Human Subjects
@@
We illustrate the brain region attention of the human subject as in Fig. 4. The first row displays the visual stimulus, while the second row illustrates the corresponding brain activities that are a 2d version of the fMRI signal. The third row shows the visual attention map, which contains similarity scores between the SOI features and each image patch. Likewise, the fourth row displays the attention map for the SOI features in relation to each patch of brain activities. The final row is the attention map generated by GradCam [81].
@@
Our analysis yielded two principal observations. Firstly, the attention map of the visual SOI features prominently focuses on the human figure within the visual stimulus. The visual attention map shares the same role as the fMRI attention map because they are generated by the same algorithm. Therefore, the fMRI attention map also shows the right ROI potentially. Secondly, detected ROIs show a consistent form across different visual stimuli. In particular, most attention scores are highly activated in OFA, PPA, OFA, and FFA regions that respond to the human face and body processing.
7.2 Non-Human Subjects
@@
We illustrate the average attention maps generated by our method and GradCam of non-human subjects as in Fig. 5. The key finding is that: Our approach generates different attention maps of different subjects, making reasonable insights for the neuroscientists. Indeed, the GradCam [81] generates similar attention maps for car, cat, dog, and truck subjects, while it also could not determine ROIs for bottles.
@@ 8 Conclusion, Limitations and Broader Impacts
This paper introduces the new BRACTIVE framework to automatically localize the visual brain region of interest. The experimental results demonstrate that the ROIs w.r.t human subjects found by BRACTIVE are highly correlated with the pre-defined ROIs from neuroscience. In addition, BRACTIVE shows its superiority by determining ROIs from various non-human subjects. This ability provides reasonable references for neuroscience fields for further discovering the human brain. Finally, BRACTIVE promises a step forward in human-like learning by leveraging brain activities to guide the DNNs, making less effort on data collection and higher performance gain.
Limitations. In our proposed method, the text is leveraged to describe the context of the visual stimulus. There is a possibility that this description does not contain the subject that the participant is focusing on. There is also a case where the participant was distracted during the experiments, making recorded fMRI signals not reflect the context of the visual stimulus.
@@
Future Work & Broader Impacts. We plan to address the limitations above and extend the BRACTIVE to different brain functionalities such as language processing, emotions, etc. More importantly, BRACTIVE promises to be a reference tool for neuroscientists and support them to debug the brain functions. With a deeper understanding of how the brain works, further novel DNNs will be inspired by these findings, boosting the AI field.
Author Biography
[Uncaptioned image] Xuan-Bac Nguyen is currently a Ph.D. student at the Department of Computer Science and Computer Engineering of the University of Arkansas. He received his M.Sc. degree in Computer Science from the Electrical and Computer Engineering Department at Chonnam National University, South Korea, in 2020. He received his B.Sc. degree in Electronics and Telecommunications from the University of Engineering and Technology, VNU, in 2015. In 2016, he was a software engineer in Yokohama, Japan. His research interests include Quantum Machine Learning, Face Recognition, Facial Expression, and Medical Image Processing.
[Uncaptioned image] Hojin Jang is an Assistant Professor in the Department of Brain and Cognitive Engineering at Korea University and the Director of the Human-Machine Cognition Lab. He is also a Research Affiliate in the Department of Brain and Cognitive Sciences at MIT. His research interests encompass Vision Science, Computational Cognitive Neuroscience, Human-AI Cognition, and Brain Decoding. He has been awarded a U.S. patent and the Jum Nunnally Dissertation Award. Dr. Jang is an active member of the Vision Sciences Society and has presented three oral presentations at its conferences.
[Uncaptioned image] Xin Li (Fellow, IEEE) received the BS degree with highest honors in electronic engineering and information science from University of Science and Technology of China, Hefei, in 1996, and the PhD degree in electrical engineering from Princeton University, Princeton, NJ, in 2000. He was a member of technical staff with Sharp Laboratories of America, Camas, WA from 2000 to 2002. He was a faculty member in Lane Department of Computer Science and Electrical Engineering, West Virginia University from 2003 to 2023. Currently, he is with the Department of Computer Science, University at Albany, Albany, NY 12222 USA. His research interests include image and video processing, compute vision and computational neuroscience. He was elected a Fellow of IEEE, in 2017 for his contributions to image interpolation, restoration and compression.
[Uncaptioned image] Samee U. Khan received a Ph.D. in 2007 from the University of Texas, Arlington, TX. Currently, he is Professor of Electrical & Computer Engineering at Mississippi State University (MSU) and served as Department Head from Aug. 2020 to July 2024. Before arriving at MSU, he was Cluster Lead (2016-2020) for Computer Systems Research at the National Science Foundation and the Walter B. Booth Professor at North Dakota State University. His research interests include optimization, robustness, and security of computer systems. His work has appeared in over 450 publications. He is the associate editor of IEEE Transactions on Cloud Computing and the Journal of Parallel and Distributed Computing. He is a Fellow of the IET and BCS. He is a Distinguished Member of the ACM and a Senior Member of the IEEE. He has won several awards, including the Best Paper Award (Systems Track), IEEE Cloud Summit, 2024; IEEE R3 Outstanding Engineer Award, 2024; IEEE Computer Society Distinguished Contributor Award, 2022 (inducted in the inaugural class); IEEE ComSoc Technical Committee on Big Data Best Journal Paper Award, 2019; IEEE-USA Professional Achievement Award, 2016; IEEE Golden Core Member Award, 2016; IEEE TCSC Award for Excellence in Scalable Computing Research (Middle Career Researcher), 2016; IEEE Computer Society Meritorious Service Certificate, 2016; Tapestry of Diverse Talents Award, North Dakota State University (NDSU), ND, USA, 2016; Exemplary Editor, IEEE Communications Surveys and Tutorials, IEEE Communications Society, 2014; Outstanding Summer Undergraduate Research Faculty Mentor Award, NDSU, ND, USA, 2013; Best Paper Award, IEEE Intl. Conf. on Scalable Computing and Communications (ScalCom), 2012; Sudhir Mehta Memorial International Faculty Award, NDSU, ND, USA, 2012; Best Paper Award, ACM/IEEE Intl. Conf. on Green Computing & Communications (GreenCom), 2010.
[Uncaptioned image] Pawan Sinha is a professor of vision and computational neuroscience in the Department of Brain and Cognitive Sciences at MIT. He received his undergraduate degree in computer science from the Indian Institute of Technology, New Delhi and his Masters and doctoral degrees in Artificial Intelligence from the Department of Computer Science at MIT. He has also had extended research stays at the University of California, Berkeley, Xerox Palo Alto Research Center, the Sarnoff Research Center in Princeton, and the Max-Planck Institute for Biological Cybernetics in Tübingen, Germany. Pawan’s research interests span neuroscience, artificial intelligence, machine learning, and public health. Using a combination of experimental and computational modeling techniques, research in his laboratory focuses on understanding the processes and principles of perceptual development. The lab’s experimental work on these issues involves studying healthy individuals and also those with neurological disorders such as autism. The goal is not only to derive clues regarding the nature and development of human visual skills, but also to create more powerful and robust AI systems. Pawan has served on the program committees for prominent scientific conferences on object and face recognition and is currently a member of the editorial board of ACM’s Journal of Applied Perception. He is a recipient of the Pisart Vision Award from the Lighthouse Guild International, the PECASE – the highest US Government award for young scientists, the Alfred P. Sloan Foundation Fellowship in Neuroscience, the John Merck Scholars Award for research on developmental disorders, the Jeptha and Emily Wade Award for creative research, the James McDonnell Scholar Award, the Troland Award from the National Academies, the Global Indus Technovator Award and the Distinguished Alumnus Award from IIT Delhi. Pawan’s teaching has been recognized by Departmental honors and the Dean’s Award for Advising and Teaching at MIT. Pawan indulges, occasionally, in adventure sports, art and whimsical projects. He has trekked to the base of Mt. Everest, jumped out of a plane at 15,000 feet, contributed a regular comic strip to the MIT campus newspaper and been inducted into the Guinness Book of Records for creating the world’s smallest reproduction of a printed book. His first journey to the United States involved a plane crash and a ride aboard the Concorde.
[Uncaptioned image] Khoa Luu is an Assistant Professor and the Director of the Computer Vision and Image Understanding (CVIU) Lab in the Department of Electrical Engineering and Computer Science (EECS) at the University of Arkansas (UA), Fayetteville, US. He is affiliated with the NSF MonARK Quantum Foundry. He is an Associate Editor of the IEEE Access Journal and the Multimedia Tools and Applications Journal, Springer Nature. He is also the Area Chair in CVPR 2023, CVPR 2024, NeurIPS 2024, WACV 2025, and ICLR 2025. He was the Research Project Director at the Cylab Biometrics Center at Carnegie Mellon University (CMU), USA. His research interests focus on various topics, including Quantum Machine Learning, Biometrics, Smart Health, and Precision Agriculture. He has received eight patents and three Best Paper Awards and coauthored 120+ papers in conferences, technical reports, and journals. He was a Vice-Chair of the Montreal Chapter of the IEEE Systems, Man, and Cybernetics Society in Canada from September 2009 to March 2011. He was the Technical Program Chair at the IEEE GreenTech Conference 2024 and a co-organizer and chair of the CVPR Precognition Workshop in 2019-2024, MICCAI Workshop in 2019 and 2020, and ICCV Workshop in 2021. He was a PC member of AAAI, ICPRAI in 2020, 2021, 2022. He has been an active reviewer for several top-tier conferences and journals, such as CVPR, ICCV, ECCV, NeurIPS, ICLR, FG, BTAS, IEEE-TPAMI, IEEE-TIP, IEEE Access, Journal of Pattern Recognition, Journal of Image and Vision Computing, Journal of Signal Processing, and Journal of Intelligence Review.

### v3

@@ 3.5 Subject of Interest Retrieval
As outlined in the previous section, it is infeasible to directly form $\textbf{P}_{s_{k}}$ and $\textbf{F}_{s_{k}}$. The Remark 1 and Remark 2 imply that if we can find the coefficient $\alpha_{ik}$, we can measure them directly. Fortunately, according to the Remark 3, $\mathbf{P}_{s_{k}}$, $\mathbf{F}_{s_{k}}$, and $\mathbf{T}_{s_{k}}$ are equivalent, but in the form of different modalities. Therefore, we measure the coefficient $\alpha_{ik}$ by measuring the similarity of $\mathbf{T}_{s_{k}}$ and each token. Motivated by this idea, we introduce the SOIR module to construct features of interest for subjects in both visual and fMRI modalities. We detail how this module operates for visual $\textbf{P}_{s_{k}}$ and note that the approach is similarly applicable for $\textbf{F}_{s_{k}}$.
Given a sequence of visual tokens $\mathbf{P}_{v}=\left[\mathbf{P}_{0},\dots,\mathbf{P}_{i},\dots,\mathbf{P}_{N_{p}-1}\right]$. Firstly, we measure coefficient $\alpha_{ik}=\texttt{sim}(\mathbf{P}_{i},\textbf{T}_{s_{k}})$. After that, we use the softmax function over these coefficients to normalize the feature aggregation. Consequently, we form $\textbf{P}_{s_{k}}$ as the weighted sum of patches’ features as mentioned in the Section 3.2. The visualization of the SOIR module is illustrated in Fig. 3.
@@ 3.6 Objective Loss Functions
Global Contextual Loss. This loss function maintains a similar global context sharing between visual, text, and fMRI data. In particular, let Contras be a contrastive loss function. The global context loss $\mathcal{L}_{g}$ is defined as in Eq. (7).
@@ 4 Brain Region of Interest Localization
As noted in Remark 2, the sub-signals or voxels in fMRI corresponding to a subject can be identified by evaluating the coefficient $\beta_{ik}$. This coefficient is defined as $\beta_{ik}=\texttt{sim}(\mathbf{F}_{s_{k}},\mathbf{F}i)$, representing the cosine similarity between $\mathbf{F}_{s_{k}}$ and $\mathbf{F}_{i}$. Additionally, a threshold $\gamma$ is predefined, designating that a sub-signal $f_{i}\in f$ is associated with the subject $s_{k}$ if $\beta_{ik}>\gamma$. The procedure for detecting the subject-specific fMRI signal is detailed in Algorithm 2. This algorithm highlights two principal observations. (1) Thanks to the Weighted SOI Loss, we can employ both kinds of features from the interested subjects, such as $\mathbf{T}_{s_{k}},\mathbf{P}_{s_{k}}$, as inputs of the algorithm. This flexibility enables the use of multiple approaches in studying brain activities. Either text or vision modalities can be used to detect the ROIs of the subjects. (2) The Algorithm 2 can be adapted for visual analysis by substituting $\textbf{F}_{f}$ with $\mathbf{P}_{v}$. This modification allows us to identify which parts of an image strongly correlate with predetermined regions of interest (ROIs) in the brain.
@@ 5 Implementation Details
Visual Encoder and fMRI Encoder. We employ the ViT-B-16 for both the visual and fMRI encoders. The input image size is $224\times 224$. The feature dimension of these encoders is set to 1024.
Text Encoder. We utilize the pre-trained transformer architecture from [4] to extract features from the text. The maximum length of the context is set to 77 [4] and the dimension of text features is also set to 1024. The text encoder is frozen while training. It will help to align visual and fMRI features with the text features easily.
@@
Datasets. To train BRACTIVE, we use the Natural Scenes Dataset (NSD) [88], a comprehensive compilation of responses from eight participants obtained through high-quality 7T fMRI scans. Each subject was exposed to approximately 73,000 natural scenes, forming the basis for constructing visual brain encoding models. Since the visual stimulus in this database is a subset of COCO [89], for each sample, there exist captions (text modality), the visual stimulus, and a corresponding fMRI response. For each subject, we split the data into five folds for training and validation.
Training Strategy. BRACTIVE is implemented in the PyTorch framework and trained using 16 $\times$ A100 GPUs (40G each). The learning rate is initially set to $2.5e^{-5}$ and then gradually reduced to zero under the CosineLinear policy [90]. The batch size is set to 44/GPU. The model is optimized by AdamW [91] for $30$ epochs. The training is completed within two hours.
@@ 6 Experimental Results
In this section, we firstly evaluate the accuracy of the proposed BRACTIVE in localizing ROIs w.r.t human subject in Section 6.1. Next, we evaluate the effectiveness of the method on downstream tasks in Section 6.2 and Section 6.3.
@@
Table I: The Dice Score of ROIs localization w.r.t human subjects. The AvgAtt is the average of the attention maps of human subjects across participants. The asterisk denotes our method trained with captions generated by the Vision-Language Model. The bold indicates the best results.
@@
Ours 69.10($\pm$ 0.3) 69.47($\pm$ 0.2) 69.50($\pm$ 0.1) 69.92($\pm$ 0.2) 69.56($\pm$ 0.2) 69.80($\pm$ 0.1) 69.82($\pm$ 0.3) 69.38($\pm$ 0.2) 72.80($\pm$ 0.2)
Table II: Performance of the brain response prediction. IMN1K refers to pretrained weights from Imagnet1K [93]. CLS and SOI refer to the CLS token and Subject of Interest features, respectively. The asterisk denotes our method trained with captions generated by the Vision-Language Model. The bold indicates the best results.
@@
Ours + SOI 50.52($\pm$ 0.3) 50.84($\pm$ 0.2) 51.48($\pm$ 0.2) 51.46($\pm$ 0.3) 50.88($\pm$ 0.2) 50.93($\pm$ 0.2) 51.72($\pm$ 0.3) 51.19($\pm$ 0.3)
Since the ROIs for specific subjects are limited, in the scope of this paper, we evaluate the accuracy of BRACTIVE with respect to human subjects only. In particular, we obtain the ROIs of the floc-faces and floc-bodies areas inside the brain as the ground truth and evaluate how well our algorithm 2 can localize the ROIs of the human subjects. In the NSD dataset, ground-truth ROIs are obtained by having each participant perform category functional localizer (fLoc) experiments [94]. These brain regions’ annotations and activations can vary across individuals due to anatomical and functional differences.. We utilize dice scores as the evaluation metric. The results are demonstrated in the Table I. The details and visualization of the prediction and ground truth can be found in the Section 7.1 and Fig. 5.
Figure 4: Region Of Interest of the human subject perceived by participant P1. The last column includes the mean of attention maps across all samples of all participants in the NSD. In the attention map, the brighter the colors, the higher the correlation to the subject.
@@
In the second experiment, we use the steps outlined in Algorithm 2 to localize the ROIs. Overall, we observe a notable enhancement ranging from 68% to 69% consistently across eight participants, compared to GradCam [92]. In the third experiment, we take the average of the attention maps across all participants. Surprisingly, the performance drops to 54.84 % for the GradCam [92] baseline while we observe a notable improvement of 72.57% from our method.
Table III: Results of the downstream tasks: object detection, instance segmentation on the COCO dataset, and semantic segmentation on the ADE20K dataset
  Pretrain $^{\text{box}} AP$$^{\text{segm}}$ mIoU
@@
Figure 5: ROI Mask w.r.t Human. The mask is generated by applying 0.5 of the threshold to the attention maps. The $6^{th}$ and $7^{th}$ columns are the ROI masks of our method and GradCam, respectively. The last column is the ground truth of ROIs w.r.t human.
6.2 Brain Response Prediction
This downstream task aims to predict the neural responses of the human brain to natural scenes, as observed during participant viewing sessions [95, 77]. We employ the pretrained visual encoder $\mathcal{E}_{v}$ to fine-tune this task. Our evaluation follows the protocol described in [95], using the Pearson Correlation Coefficient (PCC) as the metric. Results, as presented in Table II, compare different architectures and training strategies. The performance for these methods on eight participants varies between 41.02 and 41.95. In a second experiment, we use the pretrained visual encoder from BRACTIVE as initial weights and employ the features from the CLS token in the final linear layer, thereby enhancing performance by approximately 5% compared to the initial scenario. In a final experiment, settings remain similar to the second, but instead of using CLS token features, we use a weighted sum of SOI features. This adjustment leads to substantial performance improvements of 10% and 3% over the first and second experiments, respectively. The SOI features provide a richer context within the visual stimuli than the CLS features, explaining the observed enhancements in model performance.
6.3 Other Downstream Tasks: Detection and Segmentation
This section examines how the pre-trained vision encoder $\mathcal{E}_{v}$ can assist with other downstream tasks. Specifically, we select the most common vision tasks: object detection, instance segmentation, and semantic segmentation. For object detection and instance segmentation, we utilize the Mask-RCNN [96] architecture and the COCO [89] database for training and evaluation. For semantic segmentation, we utilize the UpperNet [97] architecture and the ADE20K [98] database. We conveniently utilize the base code from MMDetection and MMSegmentation [99]. The performance is shown in the Table III.
@@ 6.3 Other Downstream Tasks: Detection and Segmentation
Figure 7: ROI w.r.t Non-human Subjects.
7 Ablation Studies
@@
In this section, we demonstrate the detected ROIs of both human and non-human subjects and then analyze the meaning behind them. In addition, we also analyze the performance of the proposed method trained on the dense captions generated by Vision-Language Model (VLM).
7.1 Qualitative Results of Human Subjects’ ROI
@@
We illustrate the brain region attention of the human subject as in Fig. 4. The first row displays the visual stimulus, while the second row illustrates the corresponding brain activities, a 2D version of the fMRI signal. The third row shows the visual attention map, which contains similarity scores between the SOI features and each image patch. Likewise, the fourth row displays the attention map for the SOI features in relation to each patch of brain activities. The final row is the attention map generated by GradCam [92].
@@
Our analysis yielded two principal observations. Firstly, the attention map of the visual SOI features prominently focuses on the human figure within the visual stimulus. The visual attention map shares the same role as the fMRI attention map because they are generated by the same algorithm. Therefore, the fMRI attention map also potentially shows the right ROI. Secondly, detected ROIs show a consistent form across different visual stimuli. In particular, most attention scores are highly activated in the OFA, PPA, and FFA regions, which respond to human face and body processing.
In addition, we further illustrate the ROI mask of the human subject as in Fig. 5. The ROI masks are constructed by applying a threshold, i.e., 0.5, to the attention map. The last column is the ground truth mask, which shows the brain activations of the floc-bodies and floc-faces ROIs. The $6^{th}$ column is the ROI mask of our proposed method, while the $7^{th}$ column is the ROI mask of GradCam [92]. We measure the dice score between the predicted masks and the corresponding ground truth. The performance is reported in the Table I. The figure demonstrates that the ROI masks generated by our approach are much closer to the ground truth than the ROI masks produced by GradCam [92], which are mostly empty.
7.2 Qualitative Results of Non-Human Subjects’ ROI
@@
We illustrate the average attention maps generated by our method and GradCam for non-human subjects, as shown in Fig. 6. The key finding is that: our approach generates different attention maps for various subjects, providing valuable insights for neuroscientists. Indeed, the GradCam [92] generates similar attention maps for car, cat, dog, and truck subjects, while it also could not determine ROIs for bottles.
We further illustrate the ROIs of non-human subjects in Fig. 7. Two key findings emerge from our observations. First, the visual attention maps generated by our method are well-fitted to the subjects. Second, compared to GradCam, our method localizes the ROIs more efficiently. For instance, with bird and bottle subjects within the visual stimulus, GradCam appears to have failed in detecting the ROIs, whereas our method succeeds.
7.3 Caption Dependence of the SOIP Module
One potential limitation of BRACTIVE is its reliance on text captions to infer the subjects of interest through the SOIP module. The captions provided in the NSD dataset (adapted from MS-COCO) may omit objects that are present in the visual stimuli or contain ambiguous descriptions due to participant distraction. To investigate the impact of caption quality on BRACTIVE, we conducted an ablation study comparing different sources of captions.
We replaced the original NSD-provided COCO captions with automatically generated captions using BLIP-2 [100], a state-of-the-art Vision-Language Model capable of producing dense image descriptions. Unlike COCO captions, which typically mention one or two salient objects, BLIP-2 [100] captions cover a broader set of fine-grained objects and contextual details within the visual scene. BRACTIVE was trained with both types of captions under the same experimental protocol, and performance was measured on the standard ROI prediction tasks.
As shown in Table I and Table II, BRACTIVE trained with BLIP-2 [100] captions achieves competitive performance compared to the version trained with COCO captions, with marginal differences across metrics (within 0.3%). This indicates that dense captions generated by VLMs are a viable substitute when ground-truth captions are incomplete or unavailable. Moreover, the ability to generate captions for any new visual stimuli relaxes the dependence on manually annotated captions and broadens the applicability of BRACTIVE to datasets beyond NSD.
@@ 8 Conclusion, Limitations and Broader Impacts
This paper introduces the new BRACTIVE framework to automatically localize the visual brain region of interest. The experimental results demonstrate that the ROIs with respect to human subjects identified by BRACTIVE are highly correlated with the predefined ROIs from neuroscience. In addition, BRACTIVE demonstrates its superiority by identifying ROIs in various non-human subjects. This ability provides reasonable references for neuroscience fields, facilitating further discovery of the human brain. Finally, BRACTIVE promises a step forward in human-like learning by leveraging brain activities to guide the DNNs, requiring less effort in data collection and yielding higher performance gains.
Limitations. In our proposed method, text is leveraged to provide contextual descriptions of the visual stimulus. However, these captions may not always include the subject that the participant is focusing on, or may misalign with the brain responses in cases where the participant was distracted during the experiment. To examine this limitation, we conducted an ablation study (Section 7.3) where captions were replaced by dense descriptions automatically generated with a vision-language model. The results showed that such automatically generated captions can serve as a competitive alternative, thereby partially mitigating the reliance on dataset-provided text. This simple solution highlights the potential of reducing caption dependence through automated or multimodal captioning strategies. Nevertheless, a more ambitious direction is to reduce reliance on text altogether by developing methods that rely purely on visual and neural signals. We leave this as an interesting direction for future work.
@@
Future Work & Broader Impacts. We plan to address the limitations above and extend the BRACTIVE to different brain functionalities such as language processing, emotions, abstract art, dynamic video stimuli, or larger participant cohorts. More importantly, BRACTIVE promises to be a reference tool for neuroscientists, supporting them to debug brain functions. We see this as a promising direction for future collaborations with cognitive neuroscience labs. In addition, with a deeper understanding of how the brain works, further novel DNNs will be inspired by these findings, boosting the AI field.
9 Acknowledgment
This research is supported by the Arkansas High Performance Computing Center, which is funded through multiple National Science Foundation grants and the Arkansas Economic Development Commission.